Two metallic spheres P and Q made of same material have same smoothness but the weight of P is 8 times that of Q. If the two are heated to same temperature and left to cool then the ratio of rate of cooling of Q to that of P is
ans 2:1
1
Expert's answer
2020-03-19T11:30:38-0400

The rate at which energy radiates from the object is


"\frac{\Delta Q}{\Delta t}=e\sigma AT^4"

"\u0394Q=mc\u0394T"

"\frac{\Delta T}{\Delta t}=\frac{e\sigma AT^4}{mc}"

"m=\frac{4\pi\rho r^3}{3}\to A=4\pi r^2=4\pi\left(\frac{3m}{4\pi\rho}\right)^{\frac{2}{3}}"

"\frac{\Delta T}{\Delta t}=K\left(\frac{1}{m}\right)^{\frac{1}{3}}"

The ratio of rate of cooling of Q to that of P is  


"\left(\frac{8}{1}\right)^{\frac{1}{3}}=2"